# Towards Efficient Training of Graph Neural Networks: A Multiscale Approach

### v3

@@ Appendix F Additional Results for Shapenet Dataset
Table 18: Average epoch time for Shapenet ’Airplane’ dataset using DGCNN, using multiscale training. Time is calculated in $mili-seconds$.
@@
Table 19: Average epoch time for Shapenet ’Airplane’ dataset using DGCNN, using multiscale gradient computation training. Time is calculated in $mili-seconds$.
@@
Dataset ShapeNet Chang et al. (2015), ’Airplanes’ category
Network architecture DGCNN Wang et al. (2018) with 3 layers and 64 hidden channels
@@
Optimizer Adam Kingma & Ba (2014)
@@
Appendix G Details and Results for PPI Dataset
The PPI dataset Zitnik & Leskovec (2017) includes 20 graphs for training, 2 for validation, and 2 for testing, with an average of 2,372 nodes per graph. Each node has 50 features and is assigned to one or more of 121 possible labels. We use the preprocessed data from Hamilton et al. (2017) and present our results in Table 21. In this experiment, we use the Graph Convolutional Network with Initial Residual and Identity Mapping (GCNII) Chen et al. (2020). The details of the network architecture and the experiment are available in Table 22. Our results demonstrate that the Multiscale Gradients Computation method achieves comparable accuracy across various pooling techniques while significantly reducing training costs.
@@ Appendix G Details and Results for PPI Dataset
Dataset PPI Zitnik & Leskovec (2017)
Network architecture GCNII Chen et al. (2020) with 9 layers and 2048 hidden channels
@@
Optimizer Adam Kingma & Ba (2014)
@@
Wight initialization Xavier Glorot & Bengio (2010)

### v5

@@ Appendix H Additional Results for Shapenet Dataset
Table 21: Average epoch time for Shapenet ’Airplane’ dataset using DGCNN, using multiscale training. Time is calculated in milisecondsmili-seconds.
@@
Table 22: Average epoch time for Shapenet ’Airplane’ dataset using DGCNN, using multiscale gradient computation training. Time is calculated in milisecondsmili-seconds.
@@
Dataset ShapeNet shapenet, ’Airplanes’ category
Network architecture DGCNN dgcnn with 3 layers and 64 hidden channels
@@
Optimizer Adam kingma2014adam
@@
Appendix I Additional Inductive Learning Datasets
The PPI dataset ppi_dataset, is a biological graph dataset where nodes represent proteins, edges indicate interactions between them, and each protein is associated with multiple labels describing its biological functions. includes 20 graphs for training, 2 for validation, and 2 for testing, with an average of 2,372 nodes per graph. Each node has 50 features and is assigned to one or more of 121 possible labels. We use the preprocessed data from hamilton2017inductive and present our results in Table 24. In this experiment, we use the Graph Convolutional Network with Initial Residual and Identity Mapping (GCNII) gcnii. The details of the network architecture and the experiment are available in Table 25. Our results demonstrate that the Multiscale Gradients Computation method achieves comparable accuracy across various pooling techniques while significantly reducing training costs.
We further evaluate our methods on the NCI1 (wale2008comparison) and ogbg-molhiv (ogbn_arxiv) datasets, which are inductive learning benchmarks consisting of multiple graphs (4,110 for NCI1 and 41,127 for ogbg-molhiv). NCI1 consists of chemical compounds represented as graphs, where nodes are atoms and edges are chemical bonds, with the goal of predicting whether a compound is active against specific cancer cell lines, while MolHIV is a molecular property prediction dataset, where the task is to predict whether a molecule inhibits HIV replication based on its molecular structure. The task for both datasets is graph classification, and we note that as stated in coupette2025no those datasets are considered to be informative datasets. Our results, presented in Table˜26 demonstrate that on these multi-graph datasets and for this type of task, our methods achieve performance that is comparable to, and in some cases surpasses, the baseline results.
@@ Appendix I Additional Inductive Learning Datasets
Dataset PPI ppi_dataset
Network architecture GCNII gcnii with 9 layers and 2048 hidden channels
@@
Optimizer Adam kingma2014adam
@@
Wight initialization Xavier xavier_init
@@
Table 26: Results on NCI1 and MolHIV datasets using GCN and multiscale training with 2 levels.
Method NCI1 (Acc.) MolHIV (AUC-ROC)
Fine Grid 69.37% (0.42%) 70.39% (0.76%)
Random 70.45% (2.14%) 70.41% (1.52%)
Topk 69.37% (0.83%) 70.40% (1.62%)
Subgraph 69.73% (1.70%) 70.94% (1.59%)
Rand. & Sub. 68.56% (0.88%) 71.01% (1.35%)
Appendix J Ablation Study
We conduct an ablation study on the PubMed dataset cora_dataset. In Table˜27, we evaluate the impact of different coarsening ratios $m$ between levels. The results show that performance remains relatively stable across a range of values. For our main experiments, we choose $m=\frac{1}{2}$, as it balances two trade-offs: minimal coarsening may not yield significant training efficiency gains, while excessive coarsening—especially with multiple levels—can degrade performance.
In Table˜28, we analyze how the allocation of training epochs across levels affects results. We compare our default strategy, which doubles the number of epochs at each coarsening level, with two alternatives: using a constant number of epochs per level, and adding a fixed number of epochs at each level. All strategies yield similar accuracy, but we adopt the exponential schedule, as it emphasizes learning on coarser levels, where training is more efficient.
Table 27: Various coarsening ratios for using PubMed dataset with GCN and 3 levels of coarsening.
$m=\frac{1}{4}$ $m=\frac{1}{3}$ $m=\frac{1}{2}$ $m=\frac{4}{3}$ $m=\frac{3}{2}$
Random 77.8% (0.4%) 77.7% (0.2%) 77.5% (0.6%) 77.3% (0.5%) 77.4% (0.3%)
Topk 76.4% (0.5%) 77.5% (0.8%) 78.2% (0.3%) 76.9% (0.6%) 77.5% (0.8%)
Subgraphs 78.0% (0.3%) 77.9% (0.5%) 78.0% (0.3%) 78.0% (0.7%) 77.8% (0.4%)
Rand. & Sub. 78.2% (0.3%) 78.0% (0.5%) 77.8% (0.8%) 78.1% (0.3%) 77.7% (0.4%)
Table 28: Various epoch distributions using the PubMed dataset with GCN and 3 levels of coarsening.
Default Constant Additive
Random 77.5% (0.6%) 77.20% (0.3%) 78.10% (0.6%)
Topk 78.2% (0.3%) 76.80% (0.4%) 76.70% (0.3%)
Subgraphs 78.0% (0.3%) 78.50% (0.8%) 77.60% (0.4%)
Rand. & Sub. 77.8% (0.8%) 77.40% (0.3%) 78.00% (0.6%)